wiki:UC20181203

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 12/03/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Misty Bentz (GSU), John Bally (CU), Chip Kobulnicky (UWy), Eric Bellm (UW), Ben Williams (UW), Kal Kadlec (UW), Joanne Hughes (Seattle U), Moire Prescott (NMSU), Bill Ketzeback (APO), Mary Beth Kaiser (JHU), Sarah Tuttle (UW)


User feedback and comments from institutional representatives:

  • Colorado - users had questions about the status of DIS (discussed further below)
  • Georgia State - nothing to report
  • JHU - nothing to report
  • NAPG - no report
  • NMSU - nothing to report
  • Oklahoma - nothing to report
  • Seattle - nothing to report
  • UVa - nothing to report
  • Washington - nothing to report
  • Wyoming - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 10/24/18 through 12/01/18

0) Overview

Minimal telescope issues this period with excellent weather except for an unusual cold snap during the middle of November. The extreme cold weather (< -10 C) contributed to problems with the ARCTIC shutter and issues with the secondary mirror actuators. We also started training a new Observing Specialist, David DeColibus, an NMSU graduate student who will be acting as an emergency backup person on the 3.5-m.

1) Telescope

There were minimal telescope issues this period, the most serious of which were sticking of the Secondary Mirror Actuators A and B. The problem appeared to be cold weather related and we first slowed down the speed of the stepper motors, which worked as a stop gap, and then we replaced the axial motors with lower temperature rated devices with better torque characteristics. Unfortunately nearly a full night was lost to this problem.

2) Instruments

DIS - the blue and red cameras have been minimally operational during this period. After the October servicing the contamination on the red camera steadily increased until we needed to bring the instrument down again for servicing. The blue camera scattered light was the worst in memory and we are currently finishing up a vacuum servicing of it. During the October servicing it was identified that one possible problem on both cameras was the vacuum valves themselves. Rebuild kits were ordered and have arrived but not in time to bring each camera to atmosphere for that work. We will schedule that servicing at a later time.

TripleSpec is operational.

Agile is operational.

Echelle is operational although there was a short inadvertent warming of the instrument during this time due to operator error. This helped give a small boost to decreasing the scattered light. Inter-order light is slowly improving as expected during the colder months.

ARCTIC is operational but during the coldest nights in November we exceeded the temperature specification of the Bonn Shutter and it was sticky. We are experimenting further with the velocity and acceleration terms to improve cold weather performance.

NIC-FPS is operational.

3) ARCSAT

Has been performing nominally. We are still troubleshooting an intermittent problem with camera USB communications that has been an annoyance. During the past month a diffuser for ARCSAT was used for the first time.


Additional telescope and instrument discussion:

We've had some very cold weather early in the season and again last night, which was the first time we saw an issue with the ARCTIC shutter. Russet did some tests re: slowing the shutter speed down, which seems to help in cold weather. There is some concern about photometric response when using this slower shutter speed, so Russet and Bill are looking at reducing it to half of what it was. Russet requested that the Users Committee reps share this development re: the ARCTIC shutter speed with their users and let us know if anyone notices any reduction in photometric accuracy.

We had another telescope issue concerning the secondary actuators, which was also cold related. The engineering staff found better cold-rated motors that are also better matched to our torque requirements, and they were installed the Monday before Thanksgiving. Hopefully this problem is now behind us.

The final cold-related problem was with the tertiary mirror. Over the past two nights this required a re-map of the tertiary rotation mechanism. This is not a persistent problem per se but the port positions (i.e. the encoder for tertiary rotation) tend to drift when there are dramatic temperature changes.

We are training an NMSU Astronomy graduate student, David DeColibus, as a backup observing specialist. We had this program in place for a number of years but haven’t had one for a while; the idea is that this person would fill in if an obs-spec is sick or has some kind of emergency and cannot work a regularly scheduled shift. In order for this to work we need to keep the person current, which means scheduling them for a few nights per month, so 3.5m users will be seeing more of David as time goes on.

The continuing problems with DIS are being addressed. We ordered replacement vacuum valves; these came in but we need to bring the instrument up to atmosphere to install the replacements and then pump it down again. We also held a PDR for a DIS getter; we will implement both of these remedies as soon as the getter holding brackets are manufactured and ready to be installed (likely in Q1 2019). Until then we will service DIS if we have to/can but it will be pumped down like we’ve been doing. The current state of DIS since the last servicing is that the blue scattered light, which was VERY bad in November, is much improved. We are starting to see more scattered light on the red side. Misty asked when in Q1 will DIS be taken down and worked on, as this may impact her program. Bill and Russet responded that the servicing will be scheduled for bright time.


Q4 3.5-m scheduling:

The Q4 schedule is complete and online. Just about all of the OPEN/DD time has been scheduled; there may be 1-2 OPEN half nights remaining.


Q4 ARCSAT scheduling:

There was one week of open time remaining at the end of December, but that is now taken.


Q1 3.5-m and ARCSAT scheduling:

Ben has received all institutional requests except one. Russet is working on the schedule now and it should be out early next week. After that we will put out the ARCSAT call for Q1 proposals.


NA2 Baffle Project:

Nancy received an email from John Wilson regarding the current project status and summarized it for the group. He has identified a design for the NA2 baffle that includes the addition of two sets of 5 baffles to block rays that can reflect off the insides of the NA2 baffle and be seen at the telescope focal plane. This design seems to be effective. A comparison of the throughput at the guider detector suggests that the new design gives the same or better throughput for all locations on the guider compared to the old baffle design. The next step will be to contact Breault and Photon Engineering to solicit proposals for a professional stray light analysis.


KOSMOS Update:

The larger lead time items have been ordered and discussions are underway for the procurement of a new larger dewar to increase the hold time of the cryogen. We are still planning to have it commissioned in early Q2. Mark asked if the UW engineering team has identified which slit arrangement will work and whether the users will be expected to do nodding along the slit. He suggested that when we get a little closer to commissioning we send something out to the Users Committee or the 35m-general list that goes through a typical exposure sequence as far as what one would have to do with the instrument from acquisition to guiding to getting a full spectrum. Several people volunteered to help with this pre-commissioning activity: Eric Bellm & Melissa Graham (UW), Paula Szkody (was suggested as a helper but was not on the call to volunteer), and Chip Kobulnicky (UWy). Sarah said that prior to commissioning they plan to put together a cookbook for how the instrument works, what calibration exposures are needed, etc. They are working on nailing down these procedures now while KOSMOS is in the lab. They are also discussing the development of tools for quick-look data examination. The current pacing item is the dewar; after we have a delivery date for that we can request volunteers to help with commissioning. Eric asked what is envisioned post-commissioning in terms of instrument support and whether the instrument will support remote observing early on. Sarah said that Conor is digging into the software now and the expectation is that KOSMOS will operate within TUI like everything else. Users Committee reps are asked to spread the word that if anyone knows of previous K/COSMOS users at KPNO or Chile, please reach out to them to get involved as we move closer to getting ready to commission KOSMOS on the 3.5m.


Training and Education Working Group Update

Nancy expects to receive the final report from the TEWG in about a month.


Response to Call for White Papers for 2020 Decadal Survey

The deadline for Astro2020 white papers has been moved to February 19, 2019. Nancy would still like to prepare something for this white paper call dealing with the ground-based followup of transient sources from TESS, ZTF, LSST, gravitational wave discoveries, etc. Transient followups fall into multiple categories of the science white paper topics, and hence could be neglected because they transcend the topical boundaries. However, she needs help with this, given that this topic falls outside of her area of scientific expertise. John Bally suggested that we talk to Zach Berta-Thompson since he’s on the TESS team. Misty and Sarah agreed to help out. Someone suggested that we also request input from Jon Holtzman, the Telescope Scientist.


Debrief from ARC Board of Governors Meeting

Nancy requested permission from the ARC Board to distribute some of her slides to the Users Committee reps and/or the broader 3.5m user community. No one seemed to have a problem with that, so she will disseminate them. Mark asked whether the UC reps had received any feedback from their institutional ARC Board reps regarding the Board meeting (content, format, etc.). The only feedback (received after this telecon) was that the posting of the science highlights as supplementary material was well-received because this left more time for important discussions.


Discussion of 25th Anniversary / Strategic Planning meeting in 2019

Nancy is planning a Strategic Planning meeting in 2019 to discuss the future priorities for the 3.5m user community. This could be held in conjunction with a 25th anniversary celebration (the telescope came online in 1994!), or it doesn't have to be. Nancy proposed several options such as hold the meeting in conjunction with the summer AAS meeting (in St Louis in 2019), hold it at APO/NSO, hold it at NMSU, hold it at some other central location, etc. It will likely be in the summer. Some initial feedback from those on the call was:

  • John Bally: great idea, have it on site (at APO/NSO)
  • Joanne: having a summer meeting w/planning meeting would get a lot of people to come, but don't hold it during the summer AAS, which is usually the last week of classes for those on quarter system
  • Misty: summer is better, she tends not to attend AAS
  • Mary Beth: it would be nice for attendees to be able to see the facility if they haven't been for a while

Nancy will poll the broader user community in order to get this nailed down.


ACTION ITEMS:

Open action items from previous meetings:

  • Russet will update the calibration section of the ARCTIC users manual once the Q3/Q4 scheduling tasks are completed. STATUS: OPEN. But this is now in action.
  • UC reps should remind their users who were displaced by DIS unavailability that there is some OPEN time remaining in the Q4 schedule that they can ask for. STATUS: OPEN. A couple of open B halfs near holidays remain available.
  • Nancy will send out a draft of a white paper for the 2020 Decadal for UC feedback prior to the next telecon. STATUS: OPEN.

New action items from this meeting:

  • Users Committee reps should share the development re: the ARCTIC shutter speed with their users and let us know if anyone notices any reduction in photometric accuracy. STATUS: OPEN.
  • Users Committee reps are asked to spread the word that if anyone knows of previous K/COSMOS users at KPNO or Chile, please reach out to them to get involved as we get closer to commissioning KOSMOS on the 3.5m. STATUS: OPEN.
  • Nancy will disseminate portions of her slides from the ARC Board meeting. STATUS: OPEN.
  • Nancy will poll users on where/when to hold the 2019 strategic planning meeting. STATUS: OPEN.
  • Nancy will poll Users Committee reps on what day/time to hold our telecons in the Spring 2019 semester. STATUS: OPEN.

Next meeting:

The next Users Committee telecon will be on TBD.

Note: See TracWiki for help on using the wiki.